Tyler has a stack of cards. he picks a card records the color and returns the card to the stack. he repeats this 60 times and choose a red card 24 times. what is the experimental probability of choosing a red card from the stack
Answer by stanbon(75887)

Tyler has a stack of cards. he picks a card records the color and returns the card to the stack. he repeats this 60 times and choose a red card 24 times. what is the experimental probability of choosing a red card from the stack
----------------------------
Emperical probability = 24/60 = 2/5
-------
Cheers,
Stan H.